An aqueous solution of 6.02 M hydrochloric acid, HCl, has a density of 1.10 g/mL.

The percent by mass of HCl in the solution is
1
Expert's answer
2016-12-01T13:28:10-0500
1 mL of the 6.02 M solution of hydrochloric acid contains 0.001*6.02 moles HCl. The molar mass of HCl is 36.46 g/mol. So, the mass of HCl is 0.001*6.02*36.46 = 0.219 g HCl; the total mass of 1 mL of solution is 1.10 g. So, the fraction of HCl is 0.219/1.10 = 0.20 (or 20%).